# Minimum Weighted Feedback Arc Sets for Ranking from Pairwise Comparisons

### v2

@@ 4 Conclusion and Future Research
In this paper, we explored the relationship between the ranking problem and the Minimum Weighted Feedback Arc Set (MWFAS) problem. We demonstrated how ranking can be effectively modeled as an MWFAS problem, using an approach that balances solution quality and runtime efficiency. Using a heuristic based on the local ratio method, our proposed algorithm was able to provide fast, scalable solutions that maintained competitive accuracy across different datasets.
Our experimental results indicate that our method not only provides a practical solution for ranking large sets of agents, but also outperforms many existing methods in terms of speed. The flexibility to adapt the black-box algorithm used for MWFAS ensures that users can balance the trade-off between performance and computational cost according to their needs.
For future research, we suggest exploring other exact or heuristic methods to solve the MWFAS problem and examining their impact on the resulting ranking quality. Furthermore, investigating improved cycle detection techniques is a possible research topic. Finding alternative tie-breaking mechanisms for ranking of vertices with no directed path between them could further enhance the effectiveness of our ranking method. These directions hold promise for developing even more robust and efficient approaches to solving the ranking problem in diverse and complex settings.
In writing this paper, ChatGPT was used to assist with generating text [27].
@@ 5 Data and Code availability
Our data and code can be found in https://github.com/SoroushVahidi/Ranking_with_MWFAS. As said before, our data is the datasets used in [20]; however, we have converted their files to .txt files that include an edges list, and the input of our program is the edge list of a graph, including edges and their weights.

### v3

@@ 4 Conclusion and Future Research
In this paper, we examined the relationship between the ranking problem and the Minimum Weighted Feedback Arc Set (MWFAS) problem, and demonstrated how ranking can be formulated and solved through combinatorial optimization. Using a heuristic based on the local ratio method, our approach provides fast and scalable rankings while maintaining competitive accuracy across diverse datasets.
Our experimental results indicate that the proposed method offers a practical solution for ranking large collections of items, and achieves runtime performance significantly faster than many existing approaches in the literature. Treating the MWFAS solver as a modular black-box component further enables users to balance solution quality and computational cost by selecting appropriate algorithms depending on their application requirements.
Several directions for future work remain open. First, exploring advanced exact or heuristic methods for MWFAS may reveal improved trade-offs between computational cost and ranking accuracy. Second, developing more efficient cycle detection techniques could enhance scalability on extremely large graphs. Third, identifying principled tie-breaking mechanisms for vertices with no directed path between them may further improve ranking quality. Finally, the design of loss functions tailored to weighted ranking, beyond those studied in prior work, represents a promising research direction. These avenues offer opportunities for advancing robust and efficient ranking methodologies in diverse and complex settings.
Acknowledgment.
Portions of the text in this manuscript were drafted with the assistance of large language models (LLMs), including ChatGPT [27]. All content was reviewed and verified by the authors.
@@ 5 Data and Code Availability
All data and source code used in this study are publicly available at: https://github.com/SoroushVahidi/Ranking_with_MWFAS.
The datasets used in our experiments originate from [20] and were converted into text files containing weighted edge lists. Our implementation takes a weighted edge list as input, where each entry specifies a directed edge and its associated weight, and produces a ranking consistent with the acyclic structure induced by edge removals.